# Gravitational wave probes of particle dark matter: a review

### v3

@@ VIII.1 Transmuted black holes in binaries
Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante et al. (2017); Bell et al. (2021), eventually becoming gravitationally bound to them over time. After repeated interactions between the WIMPs and ordinary nucleons, the WIMPs become thermalized, forming a compact core with a higher density than that of the celestial object. If the density of the particles is sufficient, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH. After formation, the celestial object could accrete surrounding DM on a given timescale, and if enough matter becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object could collapse into what is known as a transmuted black hole (TBH) Dasgupta et al. (2021); Ray (2023); Bhattacharya et al. (2023, 2024); Bhattacharya (2024). If the sum of the time to accrete new particles and the time to collapse is smaller than the age of the universe, TBHs should form.
@@
TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al. (2021). Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no difference between electromagnetically silent neutron stars and low-mass BHs. Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al. (2023); Bhattacharya (2024). In LABEL:fig:wimp-cbc-cons, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al. (2023). Additionally, a projection for this constraint is shown for 50x more exposure time, which may be possible for the future GW network [^6] https://observing.docs.ligo.org/plan/.While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437--4715 has not actually collapsed into a BH in its lifetime.
@@
In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al. (2024). In LABEL:fig:lisa-wimp-cons, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors. These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime. The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band. Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form to each other.
@@
Comparing LABEL:fig:wimp-cbc-cons and LABEL:fig:lisa-wimp-cons, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM. For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al. (2023, 2024) for more details.
In contrast to TBHs, we note that if DM collects inside neutron stars, a different signature would be present. Depending on what equation-of-state assumed for baryonic or fermionic DM, neutron stars could actually be DM admixed neutron stars, which would have a particular equation-of-state that differs from the (currently unknown) canonical one. Evidence for such an effect could be found in changes to the tidal deformability parameter that could be measured by GW detectors in the future Collier et al. (2022).
VIII.2 GWs to probe the GeV excess
@@
GWs can be used to weigh in on this debate. If the millisecond pulsars are deformed, they will emit CWs as they rotate, and this radiation could be detectable by the LIGO, Virgo and KAGRA if the degree of deformation, called the ellipticity (defined in terms of the principle moments of inertia as $\epsilon=\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$), is large enough Haskell et al. (2015).
@@
First, a stochastic GW background of GWs from millisecond pulsars in the Galactic Center was considered, and projected constraints on the average ellipticity of this population were obtained as a function of the total number of millisecond pulsars in the Galactic Center Calore et al. (2019). The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in LABEL:fig:Ntoteps. However, they did not go so far as to link GW observations with Fermi-LAT ones.
On the other hand, Ref. Miller and Zhao (2023) considered different models for how the deformations on the millisecond pulsars were sustained, either through the internal magnetic field Lander et al. (2012); Mastrano and Melatos (2012); Lander (2014) or simply as a fraction of the total electromagnetic radiation from the star. It is then assumed that the unresolved population of millisecond pulsars follows the same spin and ellipticity distribution of the known pulsars Manchester et al. (2005), and, from that, the probability of observing a CW in the search be computed, given upper limits on ellipticities from searches for isolated neutron stars or those in binary systems Abbott et al. (2022b); Miller and Zhao (2023). Once this quantity is computed, we can determine the number of millisecond pulsars in the Galactic Center by fitting parameters of luminosity functions to match the observed Galactic Center GeV excess Hooper and Linden (2016); Ploeg et al. (2020); Dinsmore and Slatyer (2022). By multiplying the predicted number of millisecond pulsars with the probability of detecting them, we can rule out luminosity function parameters that would have led to at least one millisecond pulsar being detected via GWs.
@@
LABEL:fig:gev-cons shows the parameter space of a simplistic luminosity function, and the constraints that can be obtained through null results from CW searches, assuming particular ellipticity and frequency distributions for the millisecond pulsars. We can see that GW searches probe a complementary portion of the parameter space to what Fermi-LAT does: if there are more millisecond pulsars that are dimmer electromagnetically, this means more rotational power may go into GWs, and more of them are necessary to explain the GeV excess. Thus, GW searches would have more millisecond pulsars to find, and not seeing at least one of them rules out particular luminosity function parameters. Likewise, Fermi performs better when less millisecond pulsars are more luminous.
Most recently, Ref. Bartel and Profumo (2024) simulated a population of $\sim 40000$ millisecond pulsars in the Galactic Center that could explain the Galactic Center GeV excess, and found that the GW signal could lie between [200,1400] Hz, and would have strain amplitudes of $\mathcal{O}(10^{-30}-10^{-28})$ based on the population parameters chosen (e.g. moment of inertia around $\sim 10^{38}$ $\cdot$ m2, ellipticities mostly smaller than $10^{-9}$, and Boxy vs. spherical spatial distributions of the millisecond pulsars). It is argued that such signals are too weak to be detected by current and future GW observatories; however, this conclusion depends heavily on the assumed ellipticity distribution and the number of millisecond pulsars chosen to be in the population. If millisecond pulsars appear in larger numbers than assumed here, they would each be dimmer, since the luminosity of the GeV excess is fixed. In fact, the constraints from Miller and Zhao (2023) are strongest for larger populations than assumed in Bartel and Profumo (2024). Additionally, the ellipticity distribution employed in Bartel and Profumo (2024) only reaches a maximum of $\sim 2\times 10^{-9}$, which is consistent with the minimum ellipticity of neutron stars as predicted by Woan et al. (2018), but may be too conservative. Upper limits on ellipticities at the Galactic Center from CW searches have not yet constrained ellipticities to be less than $\sim 10^{-6}$ Abbott et al. (2022e, b); Steltner et al. (2023), which is consistent with the maximum ellipticity employed in Miller and Zhao (2023) up to $10^{-6}$.
If GWs can rule out more and more of the luminosity function parameter space to explain the GeV excess, then this will give significant weight to the annihilating WIMP DM hypothesis.
IX GW probes of DM spikes
IX.1 Background
It was first recognized in Gondolo and Silk (1999) that BHs in the DM halo accreting DM adiabatically (slowly) could allow a high-density region of DM to form, called a “spike”, which could enhance the DM annihilation rate inside the spike. Subsequent studies showed that if such spikes formed around supermassive BHs at the center of galaxies, they would would likely have been disrupted during host galaxy mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005). Though, recently, upper bounds on the DM densities for supermassive BHs have been forecast with LISA Daniel et al. (2025). However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005); Kim et al. (2023), since their spikes may not have been torn apart through mergers over cosmic time. Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2024); Bertone (2024); Dosopoulou and Silk (2025).
The density of the DM spike is characterized by the following equation Eda et al. (2013):
$\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\alpha}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (108)
@@ IX.1 Background
where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius. Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$. The power-law index $\alpha$ denotes how the concentration of DM particles falls off as one moves away from the host BH. Astrophysically speaking, measurements of $\alpha$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al. (2013).
IX.2 Dephasing due to dynamical friction of IMRI systems
@@
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al. (2022). Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Rahman et al. (2024); Chowdhuri et al. (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
In the future, if matched filtering analyses are performed that do not account for the DM spike, dephasing with respect to the vacuum templates may occur. In Fig. 26, we show how the signal-to-noise ratio degrades as a function of the observation time, the initial frequency $f_{\rm ini}$, and the power-law index $\alpha$ of the DM spike). As described in Eda et al. (2013), dynamical friction affects the early stages of the inspiral, and as the system progresses toward, the DM spike becomes disrupted, reducing its impact on the GW signal. That is why in Fig. 26, lower initial frequencies start to lose signal-to-noise ratio at smaller values of $\alpha$. Furthermore, it is clear that above $\alpha\simeq 2$, templates that include dynamical friction will be required to successfully recover IMRI signals in space-based GW interferometers. This conclusion was significantly expanded in Eda et al. (2015), in which the authors demonstrated how accurately the power-law index $\alpha$ could be measured from a detection of an IMRI signal in LISA. This accuracy is quantified in terms of the power-law index itself and the masses of the two objects that are inspiraling. If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately. This is because higher values of $\alpha$ induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
Figure 26: Taken from Eda et al. (2015). The number of cycles $N_{\text{cycle}}$ spent in the GW band $f\in[f_{\text{min}},f_{\text{max}}]$ for a primary BH with mass $M_{\rm BH}=1000M_{\odot}$ and a companion of $1M_{\odot}$. At large $\alpha,$ the DM solition causes large dephasing between the injected waveform and the one that would be used to recover the signal. Thus, the number of cycles that are captured in-band drops significantly.

### v4

@@ VIII GW probes of WIMP dark matter
While much of this review has focused on probing ultralight dark matter with GW interferometers, with the notable exception of macroscopic DM discussed in Section IV, it is important to also highlight the potential for using GWs to probe more conventional forms of dark matter, such as WIMPs. Although WIMPs are typically considered on different scales and their interactions are more strongly coupled than those of ultralight dark matter, they can still leave distinctive imprints on GW signals.
In this section, we outline two possible indirect ways to search for WIMP dark matter using GWs. First, in Section VIII.1, we explore the possibility that WIMPs could induce the collapse of stellar-mass objects, both compact and not, into BHs leading to unique signatures in the GW emission from these newly formed BHs. Second, in Section VIII.2, we discuss how the non-detection of CWs from the Galactic Center could help constrain the millisecond pulsar hypothesis for the observed GeV excess, thereby reinforcing the alternative scenario in which annihilating WIMP DM is the underlying cause of this astrophysical signal.
By examining these two avenues, we can see how GW observations can provide crucial insights into the nature of WIMP DM and contribute to the broader effort to identify and characterize a diverse set of DM candidates.
@@ VIII.1 Transmuted black holes in binaries
Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante et al. (2017); Bell et al. (2021), eventually becoming gravitationally bound to them over time. After WIMPs repeatedly interact with ordinary nucleons, they become thermalized, forming a compact core with a higher density than that of the celestial object. If the density of the particles becomes sufficient inside the celestial object, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH that is extremely tiny with respect to the celestial object. After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta et al. (2021); Ray (2023); Bhattacharya et al. (2023, 2024); Bhattacharya (2024). Quantitatively, if the sum of the time to accrete new particles and the time for the WIMP core to become gravitationally unstable is smaller than the age of the universe, TBHs should form.
@@
TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al. (2021). Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no clear difference between electromagnetically silent neutron stars and low-mass BHs. Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al. (2023); Bhattacharya (2024). In Fig. 24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al. (2023). Additionally, a projection for this constraint is shown for a 50x increase in the probable spacetime volume, which is reasonable for the future GW network using ET or Cosmic Explorer. While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott et al. (2012); Garani et al. (2019).
@@
In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al. (2024). In Fig. 24, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors. These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime. The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band. Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form close to each other.
@@
Comparing Fig. 24 and Fig. 24, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM. For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al. (2023, 2024) for more details.
[…]
VIII.2 GW probes of the galactic-center GeV excess
@@
GWs can be used to weigh in on this debate. If the millisecond pulsars are deformed, they will emit CWs as they rotate. The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell et al. (2015), where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
@@
One approach to constraining the millisecond pulsar hypothesis is to consider the bulk emission of a thousands or more sources of CWs that are superimposed, creating a stochastic GW background. Ref. Calore et al. (2019) modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center. The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in Fig. 25. However, they did not go so far as to link GW observations with Fermi-LAT ones.
Ref. Miller and Zhao (2023) took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs). They explored two main scenarios for the deformation of these pulsars:
  1. 1.
    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander et al. (2012); Mastrano and Melatos (2012); Lander (2014).
  2. 2.
    Model-Agnostic Deformation: In the second, more general scenario, they did not assume a specific cause for the deformation. Instead, they modeled the ellipticity as 1% of the total electromagnetic radiation emitted by the star, making it independent of the underlying physical mechanism.
Using these two models, Ref. Miller and Zhao (2023) then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester et al. (2005). By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars Abbott et al. (2022b), they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
In parallel, the number of millisecond pulsars in the Galactic Center can be inferred by modeling the electromagnetic emission associated with the GeV excess. This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper and Linden (2016); Ploeg et al. (2020); Dinsmore and Slatyer (2022). By combining the predicted number of millisecond pulsars, derived from the luminosity functions, with the probability of detecting them via CWs, the authors were able to rule out certain luminosity function parameters. Specifically, they excluded models in which the number of millisecond pulsars would have been large enough to produce at least one detectable CW, thereby exploring the possible origins of the GeV excess.
@@
Fig. 25 shows the parameter space of a luminosity function that models the observed luminosity of the Galactic Center GeV excess as following a log-normal distribution. The color show the number of millisecond pulsars predicted by each choice of the two parameters of the luminosity function ($L_{0},\sigma_{L})$. Plotted as a light blue region is an exclusion area, which signifies that CW searches would have seen at least one millisecond pulsar, assuming the aforementioned ellipticity and frequency distributions for the millisecond pulsars in the galaxy. We can see that GW searches probe a complementary portion of the parameter space to what Fermi-LAT does: if there are more millisecond pulsars that are dimmer electromagnetically, this means more rotational power may go into GWs, and more of them are necessary to explain the GeV excess. Thus, GW searches would have more millisecond pulsars to find, and not seeing at least one of them rules out particular luminosity function parameters. Likewise, Fermi performs better when less millisecond pulsars are more luminous.
[…]
In summary, this section highlights the potential of using GWs to probe WIMP DM, focusing on two distinct approaches. First, the formation of TBHs from WIMP interactions with celestial objects can lead to unique GW signatures, such as those from inspiraling neutron stars and BHs. These signals, especially in the millihertz band, offer constraints on the DM interaction cross-section and can be used to probe the properties of both bosonic and fermionic WIMPs. Second, GW non-observations of CWs from deformed millisecond pulsars in the Galactic Center provide a complementary method to weaken the millisecond pulsar hypothesis for the GeV excess observed by the Fermi-LAT experiment, which lends strength to the annihilating WIMP hypothesis for this excess. By analyzing the number and ellipticity of unresolved millisecond pulsars, GW searches can rule out certain luminosity function parameters, helping to distinguish between WIMP and millisecond pulsars explanations for the observed excess. These two complementary methods underscore the crucial role GWs can play in advancing our understanding of WIMP DM.
IX GW probes of DM spikes around inspiraling black holes
It was first recognized in Gondolo and Silk (1999) that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or through interactions with stars in the surrounding stellar population Merritt (2004); Bertone and Merritt (2005). However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian et al. (2013); Ferrer et al. (2017); Bertone (2024), which begs the question of whether these objects could survive galactic dynamics. Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone et al. (2005), it is much less clear-cut for ultralight dark matter: multiple works Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020); Ficarra (2021); Zhang et al. (2023b); Choudhary et al. (2021); Yang et al. (2018) suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea et al. (2024). The extent to which DM spikes survive directly affects the amount of dynamical friction the binary will experience as it inspirals, thus altering the GWs from the inspiral.
In light of these recent developments regarding the plausibility of forming DM spikes around supermassive BHs, Ref. Daniel et al. (2025) has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems. These bounds provide new constraints on the DM spike formation and the nature of DM in the central regions of galaxies, offering a unique opportunity to test the high-density DM environment near supermassive BH binaries. Furthermore, Ref. Chan and Lee (2024) determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao and Silk (2005); Kim et al. (2023). Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2025); Bertone (2024); Dosopoulou and Silk (2025). These findings expand the range of environments where DM spikes might form, suggesting that IMBHs could significantly influence DM distributions.
For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda et al. (2013); Cole et al. (2023a, b):
$\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\gamma_{s}}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (111)
@@ IX GW probes of DM spikes around inspiraling black holes
where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius. Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$. The power-law index $\gamma_{s}$ denotes how the concentration of DM particles falls off as one moves away from the host BH. Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al. (2013).
We consider two scenarios where dynamical friction can influence the evolution of binary black holes: in Section IX.1, we examine how dynamical friction leads to gravitational-wave dephasing in IMRI systems; in Section IX.2, we explore how DM spikes can form around PBHs and cause dephasing in binaries with primary masses of $\mathcal{O}(1\text{--}100)M_{\odot}$ and mass ratios $\leq 10^{-3}$. Lastly, in Section IX.3, we describe how observations of binary black holes can be used to place limits on DM spikes, under the assumption that their orbits remain largely unaffected by dynamical friction.
IX.1 Dephasing due to dynamical friction of intermediate mass ratio inspirals
@@
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al. (2022); Mitra et al. (2025). Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Rahman et al. (2024); Chowdhuri et al. (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
This conclusion was significantly expanded in Eda et al. (2015), in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA. This accuracy is quantified in terms of the power-law index and the masses of the two objects that are inspiraling. If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately. This is because higher values of induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al. (2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al. (2025). As an illustration, we show in Fig. 26 the impact of an environmental effect, the DM spike, on the SNR obtained when match-filtering with a vacuum template to analyze data in which a binary black hole system has a DM spike around it. Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo and Silk (1999). Ref. Cole et al. (2023b) argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation. However, it is important to note that in LISA, we will be in a regime in which signals of this kind have enormous SNR, and that parameter estimation and distinguishing among different GW signals will be a key challenge. Thus, it is more likely that parameter estimates will be biased rather than signals being completely missed. Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi et al. (2025). In principle, CW methods could provide quick point estimates for key binary black hole parameters, which would enable the development of a reduced set of templates that incorporate DM spikes. These templates could then be used for follow-up matched filtering searches.
Figure 26: Taken from Cole et al. (2023b). Percentage SNR loss from matched filtering with a vacuum template when the injected GW signal originates from a binary black hole system surrounded by DM spikes (the injected waveform includes the effect of dynamical friction). The $x$ and $y$ axes correspond to the parameters of the DM spike profile defined in Eq. 111. The assumed observation time is one year, with component masses $m_{1}=10^{5}M_{\odot}$ and $m_{2}=10M_{\odot}$.
While Ref. Cole et al. (2023b) demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref. Zwick et al. (2024) presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources. Moreover, the authors argue about the practicality of measuring such dephasing effects, noting that the early inspiral is where the effects would be most pronounced, but this is also where the GW emission is weakest; that a constant phase offset could mimic dephasing effects ; and that matched filtering over such long timescales would have practical limitations. For thee reasons.
However, Ref. Wilcox et al. (2024) provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick et al. (2024). Additionally, they propose more model-agnostic approaches, such as periodic forces, that could leave characteristic signatures on the GW signal and be applied to study the evolution of the binary system.
These differing viewpoints are crucial when considering the overall approach to detecting and interpreting GWs from systems with DM spikes. While Zwick et al. (2024) concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox et al. (2024) offer a promising alternative for tackling the problem. In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi et al. (2025). As mentioned in Section VII.2, CW methods could provide ways of addressing the three concerns laid out in Ref. Zwick et al. (2024) by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search. This approach can also be applied in lieu of traditional matched filtering for PBHs in ET and Cosmic Explorer.
IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes